# Ctrl-Adapter: An Efficient and Versatile Framework for Adapting Diverse Controls to Any Diffusion Model

### v1

@@ 5.5 Zero-Shot Generalization on Unseen Conditions
Figure 11: Zero-shot transfer of Ctrl-Adapter trained only on depth maps to unseen conditions. See full videos on our project page.
5.6 Training Efficiency of Ctrl-Adapter
In Fig. 12, we show generated images from Ctrl-Adapter with different training hours. Ctrl-Adapter learns to adapt the SDv1.5 ControlNet features to SDXL very quickly after a few hours, whereas training a ControlNet for SDXL takes 700 A100 hours. In Fig. 13, we show the training progress of Ctrl-Adapter for video and image control. Ctrl-Adapter outperforms strong baselines in less than 10 GPU hours.
Figure 12: Training efficiency of Ctrl-Adapter on SDXL backbone. Training clock time is measured on 4 A100 80GB GPUs.
Figure 13: Training speed of Ctrl-Adapter for video (left) and image (right) control with depth map. The training GPU hours are measured with A100 80GB GPUs. For both video and image controls, Ctrl-Adapter outperforms strong baselines in less than 10 GPU hours.
5.7 Video Editing via Combining Image and Video Ctrl-Adapters
In Fig. 14, we showcase that Ctrl-Adapter can make video editing easy by cascading an image diffusion model (e.g., SDXL) and an image-to-video diffusion model (e.g., I2VGen-XL) where both models are equipped with spatial control via Ctrl-Adapters. First, we extract conditions (e.g., depth map) from the original video. Next, we create the initial frame with the image diffusion model (SDXL + Ctrl-Adapter). Lastly, we provide the newly generated initial frame and frame-wise conditions to the image-to-video diffusion model (I2VGen-XL + Ctrl-Adapter) to obtain the final edited video.
Figure 14: Video editing by combining SDXL and I2VGen-XL, where both models are equipped with spatial control via Ctrl-Adapter. First, we extract conditions (e.g., depth map) from the original video. Next, we create the initial frame with SDXL + Ctrl-Adapter. Lastly, we provide the newly generated initial frame and frame-wise conditions to I2VGen-XL + Ctrl-Adapter to obtain the final edited video.
6 Ablation Studies
In the following, we show ablation studies for the design choices of Ctrl-Adapter: different adapter architectures (Sec. 6.1) and skipping the latent to ControlNet inputs (Sec. 6.2).
@@ 6 Ablation Studies
Figure 15: Comparison of different architecture of Ctrl-Adapter for image and video control, measured with visual quality (FID) and spatial control (MSE/optical flow error) metrics. The metrics are calculated from 1000 randomly selected COCO val2017 images and 150 videos from DAVIS 2017 dataset respectively. Left: image control on SDXL backbone. Right: video control on I2VGen-XL backbone. For both plots, data points in the bottom-left are ideal. SC, TC, SA, and TA: Spatial Convolution, Temporal Convolution, Spatial Attention, and Temporal Attention. ${}^{*}N$ represents the number of blocks in each Ctrl-Adapter. The diameters of bubbles represent peak training GPU memory.
6.1 Adapter Design
@@
Table 5: Skipping latent from ControlNet inputs helps Ctrl-Adapter for (1) adaptation to backbone models with different noise scales and (2) video control with sparse frame conditions.
Method Latent $\bm{z}$ is given to ControlNet Depth
FID ($\downarrow$) Optical Flow Error ($\downarrow$)
@@
SVD [3] + Ctrl-Adapter 4.33 4.07
SVD [3] + Ctrl-Adapter 4.05 3.68
@@ 6.2 Skipping Latent from ControlNet Inputs
As described in Sec. 3.2, we find skipping the latent $z$ from ControlNet inputs can help Ctrl-Adapter to more robustly handle (1) adaption to the backbone with noise scales different from SDv1.5, such as SVD and (2) video control with sparse frame conditions. Table 5 shows that skipping the latent $z$ from ControlNet improves the visual quality (FID) and spatial control (optical flow error) metrics in both scenarios.
7 Conclusion
We propose Ctrl-Adapter, an efficient and versatile framework that adds diverse controls to any image/video diffusion model, by adapting and temporally aligning pretrained ControlNets, while the parameters of the ControlNet and the backbone diffusion model are kept frozen. Training a Ctrl-Adapter is significantly more efficient than training a ControlNet for a new backbone model. Ctrl-Adapter also provides many useful capabilities including image control, video control, video control with sparse inputs, and multi-source controls. We empirically demonstrate the usefulness of Ctrl-Adapter with comprehensive analysis. With diverse image and video diffusion backbones, training Ctrl-Adapter can matche or outperform training new ControlNets, while taking much lower computational cost. Ctrl-Adapter perform zero-shot adaption to unseen conditions and help generate videos with sparse frame conditions or multiple conditions (e.g., depth map, canny edge, human pose, and surface normal). We also provide comprehensive ablations for Ctrl-Adapter design choices and qualitative examples. We hope our work can foster future works in efficient controlled generation of videos and images.
Acknowledgments
Appendix
In this appendix, we include PyTorch implementation for inverse timestep sampling (Appendix A), Ctrl-Adapter architecture details (Appendix B), comparison of Ctrl-Adapter variants and related methods (Appendix C), additional quantitative analysis (Appendix D), and additional qualitative examples (Appendix E).
Appendix A PyTorch Implementation for Inverse Timestep Sampling
In Algorithm 1, we provide the PyTorch [1] implementation of inverse timestep sampling, described in Sec. 3.2. In the example, inverse time stamping is used to adapt to the SVD [3] backbone.
{minted}
[fontsize=]python import torch
def sample_sigma(u, loc=0., scale=1.): """Draw a noise scale (sigma) from a pre-defined distribution. Here, we use lognormal distribution used in SVD.""" return torch.distributions.Normal(loc, scale).icdf(u).exp()
def sigma_to_timestep(sigma): """Map noise scale to timestep. Here, we use the function used in SVD.""" timestep = 0.25 * sigma.log() return timestep
def inverse_timestamp_sample(): """Sample noise scales and timesteps for ControlNet and diffusion models trained with continuous noise sampler. Here, we use the setting used for SVD. """ # 1) sample u from Uniform[0,1] u = torch.rand(1) # 2) calculate sigma_svd from pre-defined log-normal distribution sigma_svd = sample_sigma(u, loc=0.7, scale=1.6) # 3) calculate timestep_svd from sigma_svd via pre-defined mapping function timestep_svd = sigma_to_timestep(sigma_svd) # 4) calculate timestep and sigma for controlnet sigma_cnet, timestep_cnet = u, int(1000 * u) return sigma_svd, timestep_svd, sigma_cnet, timestep_cnet
Algorithm 1 PyTorch Implementation for Inverse Timestep Sampling
Appendix B Ctrl-Adapter Architecture Details
Figure 16: Detailed architecture of Ctrl-Adapter blocks.
In Fig. 16, we illustrate the detailed architecture of Ctrl-Adapter blocks. See Fig. 2 for how the Ctrl-Adapter blocks are used to adapt ControlNets to image/video diffusion models. We extend the Fig. 2 with adding more detailed visualizations, including skip connections, normalization layers in each module, and the linear projection layers (i.e., FFN) in each spatial/temporal attention modules.
Appendix C Comparison of Ctrl-Adapter Variants and Related Methods
In Fig. 17, we compare the variants of Ctrl-Adapter designs (with latent is given / not given to ControlNet; see Sec. 3.2 for details) and two related methods: SparseCtrl [17] and X-Adapter [48]. Unlike Ctrl-Adapter that leverages the pretrained image ControlNets, SparseCtrl (Fig. 17 c) trains a video ControlNet with control conditions $c_{f}$ and frame masks $m$ as inputs. While X-Adapter (Fig. 17 d) needs SDv1.5 UNet as well as SDv1.5 ControlNet during training and inference, Ctrl-Adapter doesn’t need to SDv1.5 UNet at all.
Figure 17: Comparison of giving different inputs to ControlNet, where $\bm{z}_{t},\bm{c}_{\text{f}}$, and $t$ represent latents, input control features, and timesteps respectively. (a): Default Ctrl-Adapter design. (b): Variant of Ctrl-Adapter where latents $\bm{z}_{t}$ are not given to ControlNet (see Sec. 3.2 for details). (c): SparseCtrl [17] trains a video ControlNet with control conditions $c_{f}$ and frame masks $m$ as inputs. (d): X-Adapter [48] needs SDv1.5 UNet as well as SDv1.5 ControlNet during training and inference, whereas Ctrl-Adapter doesn’t need to SDv1.5 UNet at all.
Appendix D Additional Quantitative Analysis
D.1 Ctrl-Adapter Design Ablations
Where to fuse Ctrl-Adapter outputs in backbone diffusion.
We compare the integration of Ctrl-Adapter outputs at different positions of video diffusion backbone model. As illustrated in Fig. 2, we experiment with integrating Ctrl-Adapter outputs to different positions of I2VGen-XL’s UNet: middle block, output block A, output block B, output block C, and output block D. Specifically, we compared our default design (Mid + Out ABCD) with four other variants (Out ABCD, Out ABC, Out AB, and Out A) that gradually remove Ctrl-Adapters from the middle block and output blocks at positions from B to D. As shown in Fig. 18, removing the Ctrl-Adapters from the middle block and the output block D does not lead to a noticeable increase in FID or optical flow error (i.e., the performances of ‘Mid+Out ABCD’, ‘Out ABCD’, and ‘Out ABC’ are similar in both left and right plots). However, Fig. 18 (right) shows that continuing to remove Ctrl-Adapters from block C causes a significant increase in optical flow error. Therefore, we recommend users retain our Ctrl-Adapters in the mid block and output blocks A, B, and C to maintain optimal performance.
Figure 18: Comparison of inserting Ctrl-Adapter to different UNet blocks. ‘Mid’ represents the middle block, whereas ‘Out ABCD’ represents output blocks A, B, C, and D. The metrics are calculated from 150 videos from DAVIS 2017 dataset.
Number of Ctrl-Adapters in each output block position.
Figure 19: Comparison of inserting different numbers of Ctrl-Adapters from the ControlNet’s zero-convolution layers to the backbone diffusion UNet’s output blocks. We use output block D here as an example. We insert three Ctrl-Adapters to three output blocks of the same feature map size by default.
Figure 20: Comparison of inserting different numbers of Ctrl-Adapters to each UNet output block. The metrics are calculated from 150 videos from DAVIS 2017 dataset. We insert 3 Ctrl-Adapters to each output block by default.
D.2 Trade-off between Visual Quality and Spatial Control

### v2

@@ 5.5 Zero-Shot Generalization on Unseen Conditions
Figure 9: Zero-shot transfer of Ctrl-Adapter trained only on depth maps to unseen conditions.
6 Downstream Tasks Beyond Spatial Control
Here, we aim to qualitatively explore how other types of ControlNets can be seamlessly integrated into our framework to enable a wide variety of downstream tasks beyond spatial control. As shown in Fig. 10, we can achieve video editing by combining image and video Ctrl-Adapters with user edited prompts; video style transfer via shuffle ControlNet $+$ Ctrl-Adapter; and text-guided motion control for masked object via inpainting ControlNet $+$ Ctrl-Adapter. See Sec. H.4 for details.
Figure 10: Illustration of additional downstream tasks supported by our Ctrl-Adapter framework.
7 Conclusion
We propose Ctrl-Adapter, an efficient, powerful, and versatile framework that adds diverse controls to any image/video diffusion model. Training an Ctrl-Adapter is significantly more efficient than training a ControlNet for a new backbone, and it can outperform or match strong baselines in visual quality and spatial control. Ctrl-Adapter not only provides many useful capabilities including image/video control, sparse frame control, multi-condition control, and zero-shot adaption to unseen conditions, but also can be easily and flexibly integrated into a variety of downstream tasks.
Acknowledgments
Appendix
[…]
Appendix A Background
A.1 Extended Related Works
Text-to-video and image-to-video generation models.
Adding control to image/video diffusion models.
While recent image/video diffusion models demonstrate impressive performance in generating highly realistic images/videos from text descriptions, it is hard to describe every detail of images/videos only with text or first frame image. Instead, there have been many works using different types of additional inputs to control the image/video diffusion models, such as bounding boxes [39, 83], reference object image [63, 17, 36], segmentation map [16, 2, 86], sketch [86], etc., and combinations of multiple conditions [33, 54, 92, 77]. As finetuning all the parameters of such image/video diffusion models is computationally expensive, several methods, such as ControlNet [86], have been proposed to add conditional control capability via parameter-efficient training [86, 64, 48]. X-Adapter [56] learns an adapter module to reuse ControlNets pretrained with a smaller image diffusion model (e.g., SDv1.5) for a bigger image diffusion model (e.g., SDXL). While they focus solely on learning an adapter for image control, Ctrl-Adapter features architectural designs (e.g., temporal convolution/attention layers) for video generation as well. In addition, X-Adapter needs the smaller image diffusion model (SDv1.5) during training and inference, whereas Ctrl-Adapter doesn’t need the smaller diffusion model at all (for image/video generation), hence being more memory and computationally efficient (see Sec. B.3 for details). SparseCtrl [21] guides a video diffusion model with conditional inputs of few frames (instead of full frames), to alleviate the cost of collecting video conditions. Since SparseCtrl involves augmenting ControlNet with an additional channel for frame masks, it requires training a new variant of ControlNet from scratch. In contrast, we leverage existing image ControlNets more efficiently by propagating information through temporal layers in adapters and enabling sparse frame control via skipping the latents from ControlNet inputs (see Sec. 3.2 for details). Furthermore, compared with previous works that are specially designed for specific condition controls on a single modality (image [86, 54] or video [31, 90]), our work presents a unified and versatile framework that supports diverse controls, including image control, video control, sparse frame control, and multi-source control, with significantly lower computational costs by reusing pretrained ControlNets (e.g., Ctrl-Adapter outperforms baselines in less than 10 GPU hours). Table 1 summarizes the comparison of Ctrl-Adapter with related works.
A.2 Extended Preliminaries: LDM and ControlNet
Latent Diffusion Models.
[…]
ControlNets.
[…]
Appendix B Ctrl-Adapter Method and Architecture Details
B.1 Ctrl-Adapter Architecture Details
Figure 11: Detailed architecture of Ctrl-Adapter blocks.
In Fig. 11, we illustrate the detailed architecture of Ctrl-Adapter blocks. See Fig. 3 for how the Ctrl-Adapter blocks are used to adapt ControlNets to image/video diffusion models. Fig. 11 is an extended version of Fig. 3 (right) with more detailed visualizations, including skip connections, normalization layers in each module, and the linear projection layers (i.e., FFN) in each spatial/temporal attention modules.
B.2 PyTorch Implementation for Inverse Timestep Sampling
In Algorithm 1, we provide the PyTorch [1] implementation of inverse timestep sampling, described in Sec. 3.2. In the example, inverse time stamping adapts to the SVD [3] backbone.
During each training step, the procedure for this algorithm can be summarized as follows:
  • Sample a variable $u$ from Uniform$[0,1]$. See line 19 in function inverse_timestamp_sample.
  • Sample noise scale $\sigma_{\text{cont.}}$ via inverse transform sampling [13]; i.e., we derive the inverse cumulative density function of $\sigma_{\text{cont.}}$ and sample $\sigma_{\text{cont.}}$ by sampling $u$: $\sigma_{\text{cont.}}=F_{\text{cont.}}^{-1}(u)$. See function sample_sigma and line 21 in function inverse_timestamp_sample.
  • Given a preconditioning function $g_{\text{cont.}}$ that maps noise scale to timestep (typically associated with the continuous-time noise sampler), we can compute $t_{\text{cont.}}=g_{\text{cont.}}(\sigma_{\text{cont.}})$. See function sigma_to_timestep and line 23 in inverse_timestamp_sample.
  • Set the timesteps and noise scales for both ControlNet and our Ctrl-Adapter as $t_{\text{CNet}}=\text{round}(1000u)$ and $\sigma_{\text{CNet}}=u$ respectively, where 1000 represents the denoising timestep range over which the ControlNet is trained. See line 25 in inverse_timestamp_sample.
During inference, we follow the similar sampling strategy, with the only change in the first step. Instead of uniformly sample a single value for $u$, we uniformly sample $k$ equidistant values for $u$ within $[0,1]$ and derive corresponding $t_{\text{cont./CNet}}$ and $\sigma_{\text{cont./CNet}}$ as inputs for denoising steps, where $k$ here is the number of denoising steps during inference.
{minted}
[linenos,fontsize=]python import torch
def sample_sigma(u, loc=0., scale=1.): """Draw a noise scale (sigma) from the noise schedule of Karras et al. (2022)""" sigma_min, sigma_max, rho = 0.002, 700, 7 # values used in the paper min_inv_rho, max_inv_rho = sigma_min ** (1 / rho), sigma_max ** (1 / rho) sigma = (max_inv_rho + (1-u) * (min_inv_rho - max_inv_rho)) ** rho return sigma
def sigma_to_timestep(sigma): """Map noise scale to timestep. Here we use the function used in SVD.""" timestep = 0.25 * sigma.log() return timestep
def inverse_timestamp_sample(): """Sample noise scales and timesteps for ControlNet and diffusion models trained with continuous noise sampler. Here we use the setting used for SVD.""" # 1) sample u from Uniform[0,1] u = torch.rand(1) # 2) calculate sigma_svd from pre-defined log-normal distribution sigma_svd = sample_sigma(u, loc=0.7, scale=1.6) # 3) calculate timestep_svd from sigma_svd via pre-defined mapping function timestep_svd = sigma_to_timestep(sigma_svd) # 4) calculate timestep and sigma for controlnet sigma_cnet, timestep_cnet = u, round(1000 * u) return sigma_svd, timestep_svd, sigma_cnet, timestep_cnet
Algorithm 1 PyTorch Implementation for Inverse Timestep Sampling
B.3 Comparison of Ctrl-Adapter Variants and Related Methods
In Fig. 12, we compare the variants of Ctrl-Adapter designs (with latent is given / not given to ControlNet; see Sec. 3.2 for details) and two related methods: SparseCtrl [21] and X-Adapter [56]. Unlike Ctrl-Adapter that leverages the pretrained image ControlNets, SparseCtrl (Fig. 12 c) trains a video ControlNet with control conditions $c_{f}$ and frame masks $m$ as inputs. While X-Adapter (Fig. 12 d) needs SDv1.5 U-Net as well as SDv1.5 ControlNet during training and inference, Ctrl-Adapter doesn’t need to SDv1.5 U-Net at all.
Figure 12: Comparison of giving different inputs to ControlNet, where $\bm{z}_{t},\bm{c}_{\text{f}}$, and $t$ represent latents, input control features, and timesteps respectively. (a): Default Ctrl-Adapter design. (b): Variant of Ctrl-Adapter where latents $\bm{z}_{t}$ are not given to ControlNet (see Sec. 3.2 for details). (c): SparseCtrl [21] trains a video ControlNet with control conditions $c_{f}$ and frame masks $m$ as inputs. (d): X-Adapter [56] needs SDv1.5 U-Net as well as SDv1.5 ControlNet during training and inference, whereas Ctrl-Adapter doesn’t need to SDv1.5 U-Net at all.
Appendix C Training and Inference Details
Model architectures.
Detailed illustration of our Ctrl-Adapter architecture has been provided across several parts of our paper, including Sec. 3, Sec. B.1, Appendix E, and Appendix F. In addition, for all the backbone models used in this paper, we kept all their parameters frozen and made no modifications.
Training details.
We use a learning rate of $5\times e^{-5}$; AdamW [44] optimizer with values for $\beta_{1}$, $\beta_{2}$, $\epsilon$, and weight decay as 0.9, 0.999, $1\times e^{-8}$, and $1\times e^{-2}$ respectively. We set the max gradient norm as 1. All our experiments are trained on 4 A100 80GB GPUs with batch size of 1. Detailed study of per-GPU training memory for different model architecture variants are shown in Fig. 13. Please note that other than mixed-precision training with data type bfloat16, we didn’t use any additional methods to speed up the training/inference clock time, or to save GPU memory. To be more specific, we didn’t use any of the following methods: xformers [35], gradient checkpointing, 8bit Adam optimizer, and DeepSpeed [59]. In addition, to make our framework easy to use directly from raw input images/videos, we extract all control condition images/frames on-the-fly during training. We train the image and video Ctrl-Adapters for 80k and 40k steps respectively, which can be finished in 24 hours measured by training clock time. The fast convergence of our method is shown in Fig. 2.
Inference details.
All inference can be done on a single A6000 GPU with 48GB memory. During inference, we use the default hyper-parameters for each backbone model, including the number of frames to generate, the number of denoising steps, and classifier-free guidance scale, etc..
Safeguards.
When we generate images during inference, we also activate the NSFW filter of the backbone models. This ensures that users are protected from unnecessary exposure to explicit or objectionable materials. For training, the datasets we used [9, 67] both filter out the image/video samples with harmful contents. For example, as stated in the “Risk mitigation” section of Panda70M paper, they used the internal automatic pipeline to filter out the video samples with harmful or violent language and texts that include drugs or hateful speech. They also use the NLTK framework to replace all people’s names with "person". LAION-POP dataset is also created by filtering out samples based on the safety tags (using a customized trained NSFW classifier that they built).
Appendix D Experimental Setup
D.1 ControlNets and Target Diffusion Models
ControlNets.
We use ControlNets trained with SDv1.5.111https://huggingface.co/lllyasviel/ControlNet SDv1.5 has the most number of publicly released ControlNets and has a much smaller training cost compared to recent image/video diffusion models. Note that unlike X-Adapter [56], Ctrl-Adapter does not need to load the source diffusion model (SDv1.5) during training or inference (see (a) and (d) in Fig. 12 for model architecture comparison).
Target diffusion models (where ControlNets are to be adapted).
For video generation models, we experiment with two text-to-video generation models – Latte [46] and Hotshot-XL [49], and two image-to-video generation models – I2VGen-XL [89] and Stable Video Diffusion (SVD) [3]. For image generation model, we experiment with PixArt-$\alpha$ [8] and the base model in SDXL [52]. For all models, we use their default settings during training and inference (e.g., number of output frames, resolution, number of denoising steps, classifier-free guidance scale, etc.).
D.2 Training Datasets for Ctrl-Adapter
Video datasets.
For training Ctrl-Adapter for video diffusion models, we download around 1.5M videos randomly sampled from the Panda-70M training set [9]. Following recent works [3, 11], we filter out videos of static scenes by removing videos whose average optical flow [14, 4] magnitude is below a certain threshold. Concretely, we use the Gunnar Farneback’s algorithm222 https://docs.opencv.org/4.x/d4/dee/tutorial_optical_flow.html [15] at 2FPS, calculate the averaged the optical flow for each video and re-scale it between 0 and 1, and filter out videos whose average optical flow error is below a threshold of 0.25. This process gives us a total of 200K remaining videos.
Image datasets.
D.3 Input Conditions
D.4 Evaluation Datasets
Video datasets.
Image datasets.
D.5 Evaluation Metrics
Visual quality.
Spatial control.
Appendix E Variants of Ctrl-Adapter Architecture Design
E.1 Ctrl-Adapter Design Ablations
E.1.1 Combinations of components within each Ctrl-Adapter
@@
Figure 13: Comparison of different architecture of Ctrl-Adapter for image and video control, measured with visual quality (FID) and spatial control (MSE/optical flow error) metrics. The metrics are calculated from 1000 randomly selected COCO val2017 images and 150 videos from DAVIS 2017 dataset respectively. Left: image control on SDXL backbone. Right: video control on I2VGen-XL backbone. For both plots, data points in the bottom-left are ideal. SC, TC, SA, and TA: Spatial Convolution, Temporal Convolution, Spatial Attention, and Temporal Attention. ${}^{*}N$ represents the number of blocks in each Ctrl-Adapter. The diameters of bubbles represent training GPU memory.
Figure 14: Comparison of inserting Ctrl-Adapter to different U-Net blocks. ‘Mid’ represents the middle block, whereas ‘Out ABCD’ represents output blocks A, B, C, and D. The metrics are calculated from 150 videos from DAVIS 2017 dataset.
Figure 15: Comparison of inserting different numbers of Ctrl-Adapters to the backbone diffusion U-Net’s output blocks. We use output block D here for illustration. We insert three Ctrl-Adapters to the output blocks of the same feature map size by default.
Figure 16: Comparison of inserting different numbers of Ctrl-Adapters to each U-Net output block. The metrics are calculated from 150 videos from DAVIS 2017 dataset. We insert 3 Ctrl-Adapters to each output block by default.
@@
E.1.2 Where to fuse Ctrl-Adapter outputs in backbone diffusion
We compare the integration of Ctrl-Adapter outputs at different positions of video diffusion backbone model. As illustrated in Fig. 3, we experiment with integrating Ctrl-Adapter outputs to different positions of I2VGen-XL’s U-Net: middle block, output block A, output block B, output block C, and output block D. Specifically, we compared our default design (Mid + Out ABCD) with four other variants (Out ABCD, Out ABC, Out AB, and Out A) that gradually remove Ctrl-Adapters from the middle block and output blocks at positions from B to D. As shown in Fig. 14, removing the Ctrl-Adapters from the middle block and the output block D does not lead to a noticeable increase in FID or optical flow error (i.e., the performances of ‘Mid+Out ABCD’, ‘Out ABCD’, and ‘Out ABC’ are similar in both left and right plots). However, Fig. 14 (right) shows that removing Ctrl-Adapters from block C causes a significant increase in optical flow error. Therefore, we recommend users retain our Ctrl-Adapters in the mid and output blocks A/B/C to ensure good performance.
E.1.3 Number of Ctrl-Adapters in each output block position
E.2 Adaptation to DiT-Based Backbones
As illustrated in Sec. 3, we have observed that the spatial features encoded in the U-Net of ControlNets and the DiT blocks are structurally different (see Fig. 22 for visualization of such observation). Therefore, mapping all middle/output blocks of ControlNet to DiT blocks might not be the optimal solution. In Fig. 17, we implement three different strategies to insert Ctrl-Adapters to the DiT blocks. Specifically, variant (a) inserts Ctrl-Adapters interleavingly into the DiT blocks, while variant (b) and (c) insert Ctrl-Adapters to the first 14 and the last 14 DiT blocks respectively. In Table 5, we perform quantitative analysis of these three variants on the DiT-based video generation model, Latte [46], with soft edge as control condition. As we can see, inserting Ctrl-Adapters interleavingly into the DiT blocks gives the best performance. This is consistent with our finding: since all DiT blocks encode global information of the generated objects, it is optimal to treat these blocks equally, rather than inserting Ctrl-Adapters only at the beginning or end. Between locations A and B, we use location A as our default setting because its feature map size ($64\times 64$) directly matches the features of the DiT blocks (also $64\times 64$) without resizing.
After finalizing where to insert Ctrl-Adapters, the next question is which block(s) of the ControlNet we should create Ctrl-Adapters to map from. In Table 6, we implement several variants on the DiT-based image generation model, PixArt-$\alpha$ [8], including mapping from the block(s) at location A, location B, location C, and location D, respectively (see Fig. 3 for the definitions of these locations). As we can see, mapping from location A or location B gives the best performance. Again, this is consistent with our findings in Fig. 22, since feature maps at locations C and D are too coarse to be informative. Moreover, we implemented two additional variants: (1) combining the ControlNet features from locations A and B (i.e., Output Blocks A+B), and (2) mapping more blocks from the same location (i.e., the second and third columns in Table 6). However, neither of these approaches provides sufficient gain compared to mapping a single block from location A or B. Therefore, we use mapping one block from location A as our default setting in our main paper.
Figure 17: Visualization of different routing methods for combining multiple ControlNet outputs. We use (a) as our default setting, and show the settings (b) and (c) as ablations.
Table 5: Ablation of inserting Ctrl-Adapters to different DiT blocks in Latte [46]. Visualization of the three architecture variants (interleaved, first half, and second half) are shown in Fig. 17. We use soft edge as control condition here for evaluation.
(a) Interleave (default) (b) First Half (c) Second Half
FID ($\downarrow$) Optical Flow Error ($\downarrow$) FID ($\downarrow$) Optical Flow Error ($\downarrow$) FID ($\downarrow$) Optical Flow Error ($\downarrow$)
18.32 2.98 19.66 3.09 23.18 3.31
Table 6: Ablation study of mapping ControlNet features from different locations, and mapping different number of blocks from the same location to the DiT blocks. The best numbers in each row are bolded, and the best numbers in each column are underscored.
Insert Locations 1 Block (default) 2 Blocks 3 Blocks
FID ($\downarrow$) SSIM ($\downarrow$) FID ($\downarrow$) SSIM ($\downarrow$) FID ($\downarrow$) SSIM ($\downarrow$)
Output Block A (default) 17.90 0.6802 19.08 0.6971 19.28 0.6855
Output Block B 18.23 0.6712 18.61 0.6720 21.47 0.6549
Output Blocks A+B 17.52 0.6812 - - - -
Output Block C 22.22 0.5273 - - - -
Output Block D 34.16 0.3506 - - - -
E.3 Skipping Latent from ControlNet Inputs
Table 7: Skipping latent from ControlNet inputs helps Ctrl-Adapter for (1) adaptation to backbone models with different noise scales and (2) video control with sparse frame conditions. We evaluate SVD and I2VGen-XL on depth maps and scribbles as control conditions respectively.
Method Latent $\bm{z}$ is given to ControlNet FID ($\downarrow$) Optical Flow Error ($\downarrow$)
@@ E.3 Skipping Latent from ControlNet Inputs
SVD [3] + Ctrl-Adapter 4.48 2.77
SVD [3] + Ctrl-Adapter 3.82 2.96
@@
As described in Sec. 3.2, we find skipping the latent $z$ from ControlNet inputs can help Ctrl-Adapter to more robustly handle (1) adaption to the backbone with noise scales different from SDv1.5, such as SVD and (2) video control with sparse frame conditions. For the first scenario, we can see from Table 7 that skipping latents in SVD leads to better visual quality (lower FID), but slightly worse spatial control (higher optical flow error). This is reasonable since skipping the noisy latents can avoid introducing large noise into the ControlNet, but it also risks losing information encoded in the latents. For the second scenario, skipping latents results in both better visual quality and better spatial control, as adding dense noisy latents can make the sparse control conditions less informative.
E.4 Different weighing modules for multi-condition generation
Table 8: Comparison of global weighting methods for multi-condition video generation (see Fig. 18 for visualization of the additional weighting methods (a.2, a.3, and a.4) developed based on (a.1) unconditional global weights). The control sources are abbreviated as D (depth map), C (canny edge), N (surface normal), S (softedge), Seg (semantic segmentation map), L (line art), and P (human pose).
D+C D+P D+C+N+S D+C+N+S+Seg+L+P
FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$)
Baseline
Equal Weights 8.50 2.84 11.32 3.48 8.75 2.40 9.48 2.93
Global MoE Router
(a.1) Unconditional Global Weights 9.14 2.89 10.98 3.32 8.39 2.36 8.18 2.48
(a.2) Timestep Emb. Weights 9.41 3.51 11.13 3.35 9.51 2.78 8.17 2.45
(a.3) Text/Image Emb. Weights 8.73 3.16 11.35 3.37 7.91 2.76 8.83 2.48
(a.4) Timestep + Text/Image Emb. Weights 8.64 3.31 10.69 3.43 8.09 2.69 8.51 2.43
Patch-Level MoE Router
(b) MLP Weights 8.40 2.34 9.37 3.17 7.87 2.11 8.26 2.00
(c) Q-Former Weights 7.54 2.39 9.22 3.22 7.72 2.31 8.00 2.08
Figure 18: Visualization of different global MoE routing methods.
Table 9: Ablation of training a unified Ctrl-Adapter v.s. training an individual Ctrl-Adapter for each condition. The results are evaluated with SDXL as the backbone model.
Training Strategy Depth Map Canny Edge
FID ($\downarrow$) SSIM ($\uparrow$) FID ($\downarrow$) SSIM ($\uparrow$)
Individual Ctrl-Adapter (default) 19.26 0.8534 21.04 0.5806
Unified Ctrl-Adapter 19.95 0.8437 22.31 0.5684
For multi-condition generation described in Sec. 3.3, in addition to the simple unconditional global weights, we also experimented with learning a router module that takes additional inputs such as diffusion time steps and image/text embeddings and outputs weights for different ControlNets. Specifically, we introduce three variants based on (a.1) unconditional global weights, which are (a.2) MLP router - taking timestep as inputs; (a.3) MLP router - taking image/text embedding as inputs; and (a.4) MLP router - taking timestep and image/text embedding as inputs. The MoE router in these variants are constructed as a 3-layer MLP. We illustrate the five methods in Fig. 18.
Table 8 show that all four global weighting schemes for fusing different ControlNet outputs perform effectively, and no specific method outperforms other methods with significant margins in all settings. With no surprise, patch-level MoE router performs consistently better than global MoE router in all control settings. Testing the effectiveness of incorporating text/image/timestep embeddings to patch-level MoE routers are left for future work.
Appendix F Additional Quantitative Analysis
F.1 Train individual Ctrl-Adapter v.s. train a unified Ctrl-Adapter
In our main paper, we train Ctrl-Adapter for each control conditions. An interesting question to ask is: can we have a single and unified Ctrl-Adapter that works for all control conditions? In this part, we perform such analysis with SDXL as the backbone model on a new training strategy. Specifically, during each training step, we randomly choose one control condition from a pool of 8 control conditions, including depth map, canny edge, soft edge, normal map, semantic segmentation, line art, user scribbles, and human pose. We train this variant with the same hyper-parameter settings as the depth map or canny edge Ctrl-Adapter mentioned in our main paper.
As shown in Table 9, training a unified Ctrl-Adapter across all control conditions does suffer from a performance decrease. However, this decrease is not very significant. Therefore, from a practical point of view, if a user has computational constraints but still needs to work on multiple control conditions, training a unified Ctrl-Adapter can be a viable workaround.
F.2 Trade-off between Visual Quality and Spatial Control